1/2 mole on helium gas is contained in a container at STP. The heat energy needed in cal to double the pressure of the gas keeping the volume constant 
 

Grade:11

1 Answers

Khimraj
5 years ago
P2 = 2P, n = ½ 
P1 = P, T1 = 273K
using
P2/T2 = P1/T1
2P/T2 = P/273
T2 = 546K
T2 – T= 273K
dQ = nCvdT
= ½*4*3*273 = 1638J
Hope it clears.